Circular No. 5793

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

NOVA AQUILAE 1993 S. J. Austin and S. G. Starrfield, Arizona State University; and R. M.

Wagner, Ohio State University, report: “Optical spectrograms of Nova Aql 1993 were obtained with the Perkins 1.8-m telescope and CCD spectrograph on May 17.35 UT. The spectra (range 330-870 nm, resolution 1.2 nm) exhibit strong permitted emission lines due to the Balmer series of hydrogen; Fe II (multiplets 37, 38, 42, 48, 49, 73, and 74); O I 777.4 and 844.6 nm; Ca II (H, K, and the infrared triplet); Na D; and Mg II 823.2 nm. Many of these lines show P-Cyg profiles. The Hα emission line profile is very asymmetric to the blue side of line center. The spectra confirm the identification as a ‘Fe II class’ nova in the early post-maximum phase of its development.”

V1974 CYGNI P. Rafanelli and L. Rosino, Department of Astronomy, University of

Padua, report: “CCD spectra of the ‘neon nova’ V1974 Cyg (N Cyg 1992) obtained on Mar. 19 and Apr. 21 with the 1.82-m telescope at Ekar (Asiago) show that the nova is now in a stage of very high ionization with outstanding nebular lines of [Ne V] 334.6, 342.6 nm; [Ne III] 386.9, 396.8 nm; [Ne IV] 472.0 nm; and [O III] 436.3, 495.9, 500.7 nm. The intensities of these lines, relative to Hβ, were on Mar. 19 several times higher than Hβ (6.4 × 10^{−12} erg cm^{−2} s^{−1}), ranging from a maximum of ∼ 49 for [Ne V] at 342.6 nm to a minimum of nearly 2.0 times Hβ for [Ne IV] at 472.0 nm. Many other lines from ions of high ionization potential are due to He II 468.6 nm, [Ne II] 575.5 nm, [Fe VI], and [Fe VII] 608.7 nm. The following coronal lines have also been observed: [Fe XIV] 530.3 nm, [A X] 553.5 nm, [Fe X] 637.5 nm, and very likely [Fe XI] 789.2 nm. Nebular and coronal lines are presently strengthening relative to Hβ, while the relative intensities of permitted lines of lower ionization potential (N II, He I) are slowly fading. The line intensity ratio Hα/Hβ was nearly 3 in March and April.”

NOVA OPHIUCHI 1993 Further visual magnitude estimates (cf. IAUC 5777): Apr. 29.72 UT,

11.5 (A. F. Jones, Nelson, N.Z.); May 1.05, 11.7 (L. Szentaskó, Budapest, Hungary); 11.95, 12.9 (Szentaskó); 12.95, 13.0 (Szentaskó); 13.48, 12.6 (Jones); 15.23, 12.9 (W. G. Dillon, Missouri City, TX).

1993 May 17 Daniel W. E. Green